# Multi-Order Wavelet Derivative Transform for Deep Time Series Forecasting

### v1

@@ Discussion
Recent advances in time series forecasting largely utilize FRL-based approaches to boost efficiency by focusing on crucial time-frequency features with minimal computational overhead. While this strategy has proven effective, WaveTS introduces a fresh perspective by leveraging the WDT to capture multi-scale stationary features in a more efficient and energy-preserving manner, amplifying sudden changes in time series data. We hope that this new framework not only enriches the LTSF community’s toolbox but also encourages the emergence of robust, scalable time series forecasting solutions. We provide discussion on the limitations of WaveTS in Appendix E.
@@ 5 Conclusion and Future Work
This paper introduces WaveTS, which employs a multi-branch architecture to achieve superior results in deep time series forecasting. In contrast to previous forecasting models that utilize the Fourier transform for feature extraction, WaveTS effectively captures multi-scale time-frequency features, highlighting regime changes in time series through a theoretically sound Wavelet Derivative Transform, while preserving both information and energy. Empirical results on long-term forecasting task and short-term forecasting task demonstrate that WaveTS outperforms previous FRL-based and Transformer-based methods, concurrently maintaining low computational complexity. Future research will concentrate on optimizing the WaveTS architecture for practical deployment.
@@
[…]
Appendix A Theoretical Proofs
A.1 Wavelet Derivative Transform
Lemma A.1.
Given a time series $X(t)$, the wavelet transform of the $n$order derivative of $X(t)$, namely Wavelet Derivative Transform (WDT), is related to the wavelet transform of $X(t)$ using the $n$-th derivative of the mother wavelet $\psi^{(n)}(t)$ as follows:
@@ A.1 Wavelet Derivative Transform
[…]
Using integration by parts $n$ times, and assuming that boundary terms vanish (i.e., the boundary term $\sum_{m=0}^{n-1}(-1)^{m}\Bigl{[}X^{(n-1-m)}(t)\,\psi_{k,j}^{(m)}(t)\Bigr{]}_{-\infty}^{+\infty}$ sum to zero. In practice, this assumption is typically satisfied because many mother wavelets, such as Daubechies wavelets, are either compactly supported or decay rapidly [4, 5], making $\psi_{k,j}^{(m)}(t)$ negligible at the boundaries. We then transfer the derivative from $X^{(n)}(t)$ to the wavelet function:
[…]
@@
[…]
@@
This result establishes the theoretical relationship between the wavelet transform of the $n$-th derivative of a time series and the wavelet transform conducted with the $n$-th derivative of the mother wavelet. To the best of our knowledge, this is the first theoretical analysis devoted to wavelet derivative learning. In practice, however, one need not explicitly differentiate the mother wavelet function. Although the derivative formally relies on $\psi^{(n)}$, standard discrete wavelet filters (such as Daubechies [5]) already capture difference and derivative features at fine scales. Hence, one can simply apply the usual WT to the original signal and then scale the resulting coefficients (by factors depending on the scale and derivative order) to emulate higher-order derivatives. This avoids numerical differentiation and the complexity of directly constructing or discretizing $\psi^{(n)}$. Wavelets with sufficient vanishing moments inherently act as finite-difference operators, thereby providing a natural approximation to derivative information. The linearity and multi-scale nature of the wavelet transform make it especially attractive for derivative-related tasks in signal processing [29, 5].
Theorem A.2 (Energy Conservation of WDT).
@@
From the standard WT Parseval relation [35], we know that (up to a wavelet-dependent constant):
@@
[…]
@@
A.2 LTI State Space Models are FRL-based Forecasters
State space models (SSMs) are highly effective for time series modeling due to their inherent simplicity and capacity to capture complex autoregressive dependencies. By integrating observed data with latent state variables, SSMs can accurately delineate dynamic patterns and the underlying processes present within the data [2, 14, 13]. This paper posits that SSMs serve as FRL-based forecasters in certain contexts; therefore, the forecasting while backcasting training approach consistently exhibits superior performance when applied to SSMs, corroborating the findings presented in Table 10. Here, we present a two-stage proof demonstrating a direct correspondence between Linear Time-Invariant (LTI) SSMs and frequency-domain linear mappings utilized by FRL-based forecasters.
@@ A.2 LTI State Space Models are FRL-based Forecasters
By the Convolution Theorem [33], we have:
@@
Appendix B Experimental Details
B.1 Datasets
We evaluate WaveTS using 10 real-world datasets, with detailed information provided in Table 4 for long-term series forecasting and Table 5 for short-term series forecasting. Specifically, the “Dimension” column indicates the number of variables in each dataset. “Forecasting Length” specifies the number of future time points to forecast, with four different forecasting horizons included for each dataset. “Dataset Size” represents the total number of time points in the training, validation, and test splits. “Information (Frequency)” denotes the domain and the sampling interval of the time series. Specifically, ETT comprises seven electricity transformer variables collected from July 2016 to July 2018. We utilize four subsets: ETTh1 and ETTh2 are sampled hourly, while ETTm1 and ETTm2 have 15-minute intervals. Exchange includes daily exchange rate data from eight countries spanning 1990 to 2016. Weather consists of 21 meteorological factors recorded every 10 minutes at the Weather Station of the Max Planck Bio-geochemistry Institute in 2020. Electricity captures hourly electricity consumption data from 321 clients. Traffic contains hourly road occupancy rates measured by 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. Additionally, we incorporate the PEMS03, PEMS04, PEMS07, and PEMS08 datasets, which are traffic datasets from the California Transportation Agencies, capturing various aspects of traffic flow and congestion. The M4 dataset comprises 100,000 time series from diverse domains, including finance, economics, and demographics, providing a comprehensive benchmark for evaluating forecasting methods across multiple scales and frequencies. Datasets are provided from [47, 61]. All datasets undergo the same preprocessing steps and are split into training, validation, and test sets following the chronological order to prevent data leakage, as outlined in TimesNet [47].
@@ B.2 Baselines
FITS [49]: FITS is a lightweight model for time series analysis that operates by interpolating time series in the complex frequency domain. This approach enables accurate forecasting and anomaly detection with only about 10k parameters, making it suitable for edge devices. We download the official code from: https://github.com/VEWOXIC/FITS.
@@
DeRiTS [7]: DeRiTS addresses distribution shifts in non-stationary time series by utilizing the entire frequency spectrum through Frequency Derivative Transformation (FDT) and an Order-adaptive Fourier Convolution Network, enhancing forecasting performance and shift alleviation. Due to the absence of an open-source implementation of DeRiTS, we undertook the task of reproducing the code independently to replicate all experiments.
@@
FilterNet [53]: FilterNet enhances Transformer-based forecasting by incorporating learnable frequency filters that selectively pass or attenuate signal components. This method improves robustness to high-frequency noise and leverages the full frequency spectrum for more accurate predictions. We download the official code from: https://github.com/aikunyi/FilterNet.
@@
WPMixer [31]: In WPMixer, multi-resolution wavelet transform extracts joint time–frequency features; the resulting sub-series are patched and embedded to extend history and preserve local patterns. An MLP-Mixer then globally mixes the patches, enabling the model to exploit both local and global information. However, a careful review of its official code revealed a fatal drop last bug (see line 23 in data_factory.py in https://github.com/Secure-and-Intelligent-Systems-Lab/WPMixer). Specifically, they excluded a significant portion of test data—especially with large batch sizes (128 and 256 in their experiments), resulting in unfair comparisons. Therefore, we fixed the bug and rerun their experiments following the reported configurations.
@@
TimeMixer [41]: TimeMixer regards a time series as a set of disentangled fine-scale (seasonal) and coarse-scale (trend) signals, then fuses them bidirectionally: Past-Decomposable-Mixing (PDM) mixes seasonal and trend components from fine to coarse and coarse to fine, while Future-Multipredictor-Mixing (FMM) ensembles multiple predictors across those scales. We download the official code from: https://github.com/kwuking/TimeMixer.
@@
iTransformer [26]: iTransformer reconfigures standard Transformer models for time series forecasting by switching input dimensions and leveraging attention mechanisms across variate tokens. This adaptation improves the model’s capability to learn temporal patterns. We download the official code from: https://github.com/thuml/iTransformer.
@@
PatchTST [32]: PatchTST modifies Transformer architectures with patch-based processing and a channel-independent framework, optimizing performance for long-term forecasting tasks. Its design is particularly effective for capturing intricate temporal structures. We download the official code from: https://github.com/PatchTST.
@@
TimesNet [47]: TimesNet introduces a task-agnostic framework that reshapes one-dimensional time series into two-dimensional tensors, enabling it to address multi-periodic patterns. We download the official code from: https://github.com/thuml/Time-Series-Library.
FEDformer [63]: FEDformer incorporates a novel attention mechanism operating in the frequency domain, combining low-rank approximations and mixed decomposition techniques to reduce distribution shifts and enhance forecasting accuracy. We download the official code from: https://github.com/MAZiqing/FEDformer.
B.3 Implementation Details

### v2

@@ Discussion
Our proposed WaveTS expands the LTSF and the STSF toolbox and promotes the development of efficient, robust and scalable forecasting solutions through time-frequency transformation. We also provide discussion on the limitations of WaveTS in Appendix F.
@@ 5 Conclusion and Future Work
This paper introduces WaveTS, which employs a multi-branch architecture to achieve superior results in deep time series forecasting. In contrast to previous forecasting models that utilize the Fourier transform for feature extraction, WaveTS targets modeling the derivative of time series rather than the raw data, effectively captures multi-scale time-frequency features, highlighting regime changes in time series through a theoretically sound Wavelet Derivative Transform, while preserving both information and energy. Empirical results on long-term forecasting task and short-term forecasting task demonstrate that WaveTS outperforms previous FRL-based and Transformer-based methods, concurrently maintaining low computational complexity. Future work will aim to optimize the WaveTS architecture for real-world deployment and broaden its applicability to downstream tasks, including classification and anomaly detection.
Reproducibility statement
We have taken several steps to ensure the reproducibility of our work. The model architecture is described in Section 3.2 of the main text. The theoretical proofs of our proposed algorithm are provided in Appendix B. The dataset statistics and descriptions are given in Appendix C.1. The implementation details, including training settings and hyperparameters, are provided in Appendix C.3. Furthermore, we include an anonymous link to the source code and scripts in the Abstract, which enables independent reproduction of our results.
@@
Appendix A Frequency Domain Representation Learning
Figure 7: The paradigm of Frequency domain Representation Learning. Domain transformation is performed by the Discrete Fourier Transform (DFT) here.
We first clarify the definition of Frequency domain Representation Learning (FRL) for better understanding. Let $X_{1},X_{2},\ldots,X_{T}]\in\mathbb{R}^{T\times C}$ denote a multivariate time series, where each $X_{t}\in\mathbb{R}^{C}$ contains the observations of $C$ variates at time $t$. At time $t$, the input window is $\mathbf{X}_{t}=[X_{t-L+1},\ldots,X_{t}]\in\mathbb{R}^{L\times C}$ and the forecasting target is $\mathbf{Y}_{t}=[\hat{X}_{t+1},\ldots,\hat{X}_{t+\tau}]\in\mathbb{R}^{\tau\times C}$. FRL models this map in the frequency domain: a bijective transform $\mathcal{T}$ produces coefficients $\mathbf{Z}_{t}=\mathcal{T}(\mathbf{X}_{t})$; a learnable operator $f(\cdot\,;\theta)$ acts on the coefficients to yield $\widehat{\mathbf{Z}}_{t}=f(\mathbf{Z}_{t}\,;\theta)$; the inverse transform reconstructs the prediction $\widehat{\mathbf{Y}}_{t}=\mathcal{T}^{-1}(\widehat{\mathbf{Z}}_{t})$ (see Figure 7). In the Fourier case, the coefficients comprise real and imaginary parts, or magnitude and phase; in the wavelet case, they comprise an $LL$ trend and multilevel $\mathbf{LH}$ details. Working on coefficients separates scales and oscillations, mitigates nonstationarity, enables band aware filtering and denoising, and remains lightweight because $f$ operates on compact structured representations. The effectiveness of FRL is that the frequency spectrum exposes seasonality and other global regularities, and provides a holistic view of temporal behavior while supporting multi scale and multi frequency analysis Yi et al. (2024b). Therefore, FRL has been widely integrated across deep backbones for time series forecasting, including Transformer based designs (Zhou et al., 2022b; Eldele et al., 2024), graph neural networks (Cao et al., 2020; Yi et al., 2023a), and MLP style architectures (Yi et al., 2023b; Fan et al., 2014; Xu et al., 2024; Fei et al., 2025). In this paper, WaveTS instantiates FRL with the WDT, which is real valued with an exact inverse: we transform $\mathbf{X}_{t}$ to WDT coefficients, apply a compact branch wise mapping across derivative order and multilevel scale, then reconstruct $\widehat{\mathbf{Y}}_{t}$ by $\mathcal{T}^{-1}$.
Appendix B Theoretical Proofs
B.1 Wavelet Derivative Transform
Lemma B.1.
Given a time series $X(t)$, the wavelet transform of the nnorder derivative of $X(t)$, namely Wavelet Derivative Transform (WDT), is related to the wavelet transform of $X(t)$ using the $n$-th derivative of the mother wavelet $\psi^{(n)}(t)$ as follows:
@@ B.1 Wavelet Derivative Transform
$W_{X^{(n)}}(k,j)=\int_{-\infty}^{+\infty}X^{(n)}(t)\psi_{k,j}(t)dt.$ (11)
Using integration by parts $n$ times, and assuming that boundary terms vanish (i.e., the boundary term $\sum_{m=0}^{n-1}(-1)^{m}\Bigl[X^{(n-1-m)}(t)\,\psi_{k,j}^{(m)}(t)\Bigr]_{-\infty}^{+\infty}$ sum to zero. In practice, this assumption is typically satisfied because many mother wavelets, such as Daubechies wavelets, are either compactly supported or decay rapidly (Daubechies, 1988; 1992), making $\psi_{k,j}^{(m)}(t)$ negligible at the boundaries. We then transfer the derivative from $X^{(n)}(t)$ to the wavelet function:
$W_{X^{(n)}}(k,j)=(-1)^{n}\int_{-\infty}^{+\infty}X(t)\frac{d^{n}}{dt^{n}}\psi_{k,j}(t)dt.$ (12)
@@
$\begin{split}W_{X^{(n)}}(k,j)&=(-1)^{n}\int_{-\infty}^{+\infty}X(t)\frac{d^{n}}{dt^{n}}\psi_{k,j}(t)dt\\ &=(-1)^{n}2^{k/2+nk}\int_{-\infty}^{+\infty}X(t)\psi^{(n)}(2^{k}t-j)dt.\end{split}$ (14)
@@
This lemma establishes a precise relationship between the WT of the $n$-th derivative $X^{(n)}(t)$ and the WT of $X(t)$ computed with the derivative wavelet $\psi^{(n)}$, where the only difference is a fixed scale-dependent gain factor $2^{nk}$ (and a sign $(-1)^{n}$).
Practical implementation (no explicit $\psi^{(n)}$). Although the identity above is written in terms of $\psi^{(n)}$, we do not numerically differentiate the mother wavelet in practice. Instead, we implement WDT as one standard DWT + scale-wise rescaling: we first compute the standard DWT coefficients of $X(t)$ using $\psi$, denoted as $W_{X}(k,j)$, and then define the $n$-th order WDT coefficients by
$\mathrm{WDT}^{(n)}_{k,j}\;=\;\alpha_{n,k}\,W_{X}(k,j),\qquad\alpha_{n,k}\;=\;(-1)^{n}2^{nk}.$ (18)
In vectorized form consistent with our notation $\mathrm{WDT}_{k}^{(n)}$, the rescaling in Eq. 18 is applied element-wise to all translations $j$ at the same scale $k$.
Why rescaling is theoretically justified. Wavelets with sufficient vanishing moments act as multiscale differential operators. In particular, when $\psi$ has $n$ vanishing moments, there exists a smoothing function $\theta$ such that $\psi(t)=(-1)^{n}\tfrac{d^{n}}{dt^{n}}\theta(t)$, and the (continuous) wavelet coefficients admit the differential form (see (Mallat, 1989; Daubechies, 1992))
$W_{X}(u,s)\;=\;s^{n}\,\frac{d^{n}}{du^{n}}\bigl(X*\theta_{s}\bigr)(u),$ (19)
which implies that standard wavelet coefficients already encode $n$-th order derivative information up to a scale gain. Importantly, our discrete wavelet basis uses the convention $\psi_{k,j}(t)=2^{k/2}\psi(2^{k}t-j)$, i.e., the dilation is $2^{k}$ (rather than $2^{-k}$). Under this convention, the continuous scale parameter corresponds to $s=2^{-k}$, hence the factor $s^{-n}$ in Eq. 19 becomes $2^{nk}$, exactly matching the scale-wise gain in Eq. 18 and the $2^{nk}$ factor derived in the lemma. Therefore, Eq. 18 provides a theoretically sound realization of derivative-aware wavelet analysis while bypassing explicit construction of $\psi^{(n)}$.
Numerical validity under db1 (Haar). In our implementation, we unify the mother wavelet to db1. This choice makes the above strategy particularly rigorous because db1’s high-pass analysis filter is $G(z)=\tfrac{1}{\sqrt{2}}(1-z^{-1})$, i.e., the discrete first-difference operator (up to normalization). Consequently, higher-order behavior corresponds to the classical binomial difference operator
$(1-z^{-1})^{n}\;=\;\sum_{m=0}^{n}\binom{n}{m}(-1)^{m}z^{-m},$ (20)
which is a standard and stable numerical approximation to $\tfrac{d^{n}}{dt^{n}}$ in the discrete domain (LeVeque, 2007). Thus, with db1, “standard DWT + scale-wise rescaling” is not a heuristic workaround, but a robust discrete realization of multiscale finite differences. Moreover, this implementation remains efficient: it requires only a single standard DWT (typically $O(T)$ for length-$T$ signals) plus per-scale scalar multiplications.
Theorem B.2 (Energy Conservation of WDT).
@@
From the standard WT Parseval relation (Parseval, 1806), we know that (up to a wavelet-dependent constant):
@@
$\sum_{k,j}\bigl|W_{X^{(n)}}(k,j)\bigr|^{2}\;=\;\sum_{k,j}\bigl|(-1)^{n}\,2^{n\,k}\,\tilde{W}^{(n)}_{k,j}\bigr|^{2}\;=\;\sum_{k,j}2^{2\,n\,k}\,\bigl|\tilde{W}^{(n)}_{k,j}\bigr|^{2}.$
@@
B.2 LTI State Space Models are FRL-based Forecasters
State space models (SSMs) are highly effective for time series modeling due to their inherent simplicity and capacity to capture complex autoregressive dependencies. By integrating observed data with latent state variables, SSMs can accurately delineate dynamic patterns and the underlying processes present within the data (Behrouz et al., 2024; Hu et al., 2024b; a). This paper posits that SSMs serve as FRL-based forecasters in certain contexts; therefore, the forecasting while backcasting training approach consistently exhibits superior performance when applied to SSMs, corroborating the findings presented in Table 10. Here, we present a two-stage proof demonstrating a direct correspondence between Linear Time-Invariant (LTI) SSMs and frequency-domain linear mappings utilized by FRL-based forecasters.
@@ B.2 LTI State Space Models are FRL-based Forecasters
By the Convolution Theorem (Oppenheim et al., 1999), we have:
@@
Appendix C Experimental Details
C.1 Datasets
We evaluate WaveTS using 10 real-world datasets, with detailed information provided in Table 4 for long-term series forecasting and Table 5 for short-term series forecasting. Specifically, the “Dimension” column indicates the number of variables in each dataset. “Forecasting Length” specifies the number of future time points to forecast, with four different forecasting horizons included for each dataset. “Dataset Size” represents the total number of time points in the training, validation, and test splits. “Information (Frequency)” denotes the domain and the sampling interval of the time series. Specifically, ETT comprises seven electricity transformer variables collected from July 2016 to July 2018. We utilize four subsets: ETTh1 and ETTh2 are sampled hourly, while ETTm1 and ETTm2 have 15-minute intervals. Exchange includes daily exchange rate data from eight countries spanning 1990 to 2016. Weather consists of 21 meteorological factors recorded every 10 minutes at the Weather Station of the Max Planck Bio-geochemistry Institute in 2020. Electricity captures hourly electricity consumption data from 321 clients. Traffic contains hourly road occupancy rates measured by 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. Additionally, we incorporate the PEMS03, PEMS04, PEMS07, and PEMS08 datasets, which are traffic datasets from the California Transportation Agencies, capturing various aspects of traffic flow and congestion. The M4 dataset comprises 100,000 time series from diverse domains, including finance, economics, and demographics, providing a comprehensive benchmark for evaluating forecasting methods across multiple scales and frequencies. Datasets are provided from (Wu et al., 2023; Zhou et al., 2021). All datasets undergo the same preprocessing steps and are split into training, validation, and test sets following the chronological order to prevent data leakage, as outlined in TimesNet Wu et al. (2023).
@@ C.2 Baselines
FITS Xu et al. (2024): FITS is a lightweight model for time series analysis that operates by interpolating time series in the complex frequency domain. This approach enables accurate forecasting and anomaly detection with only about 10k parameters, making it suitable for edge devices. We download the official code from: https://github.com/VEWOXIC/FITS.
@@
DeRiTS Fan et al. (2014): DeRiTS addresses distribution shifts in non-stationary time series by utilizing the entire frequency spectrum through Frequency Derivative Transformation (FDT) and an Order-adaptive Fourier Convolution Network, enhancing forecasting performance and shift alleviation. Due to the absence of an open-source implementation of DeRiTS, we undertake the task of reproducing the code independently to replicate all experiments.
@@
TexFilter Yi et al. (2024a): TexFilter enhances Transformer-based forecasting by incorporating learnable frequency filters that selectively pass or attenuate signal components. This method improves robustness to high-frequency noise and leverages the full frequency spectrum for more accurate predictions. We download the official code from: https://github.com/aikunyi/FilterNet.
@@
WPMixer Murad et al. (2025): In WPMixer, multi‑resolution wavelet transform extracts joint time–frequency features; the resulting sub‑series are patched and embedded to extend history and preserve local patterns. An MLP‑Mixer then globally mixes the patches, enabling the model to exploit both local and global information. However, a careful review of its official code revealed a fatal drop last bug (see line 23 in data_factory.py in https://github.com/Secure-and-Intelligent-Systems-Lab/WPMixer). Specifically, they excluded a significant portion of test data—especially with large batch sizes (128 and 256 in their experiments), resulting in unfair comparisons. Therefore, we fix the bug and rerun their experiments following the reported configurations.
@@
TimeMixer Wang et al. (2024a): TimeMixer regards a time series as a set of disentangled fine‑scale (seasonal) and coarse‑scale (trend) signals, then fuses them bidirectionally: Past‑Decomposable‑Mixing (PDM) mixes seasonal and trend components from fine to coarse and coarse to fine, while Future‑Multipredictor‑Mixing (FMM) ensembles multiple predictors across those scales. We download the code from: https://github.com/kwuking/TimeMixer for short-term forecasting.
TimeMixer++ Wang et al. (2025b): TimeMixer++ is a versatile model that excels at various time series tasks by effectively capturing and extracting complex patterns across multiple time scales and frequency resolutions. Through techniques like multi-resolution time imaging and various mixing strategies, TimeMixer++ achieves state-of-the-art performance, outperforming both general-purpose and task-specific models in forecasting, classification, and other time series analyses. Although an GitHub repository implementing TimeMixer++ is available at https://anonymous.4open.science/r/TimeMixerPP , rerunning the code does not reproduce the results reported in the paper. Therefore, we adopt the results directly from the paper only for long-term forecasting task. For Table 3, however, we rerun the provided code to obtain the reported efficiency metrics.
@@
iTransformer Liu et al. (2024c): iTransformer reconfigures standard Transformer models for time series forecasting by switching input dimensions and leveraging attention mechanisms across variate tokens. This adaptation improves the model’s capability to learn temporal patterns. We download the official code from: https://github.com/thuml/iTransformer.
@@
PatchTST Nie et al. (2023): PatchTST modifies Transformer architectures with patch-based processing and a channel-independent framework, optimizing performance for long-term forecasting tasks. Its design is particularly effective for capturing intricate temporal structures. We download the official code from: https://github.com/PatchTST.
@@
TimesNet Wu et al. (2023): TimesNet introduces a task-agnostic framework that reshapes one-dimensional time series into two-dimensional tensors, enabling it to address multi-periodic patterns. We download the official code from: https://github.com/thuml/Time-Series-Library.
DLinear Zeng et al. (2023): DLinear questions the suitability of Transformers for long-term time series forecasting and proposes a simple one-layer linear model (LTSF-Linear) as a baseline. Surprisingly, LTSF-Linear consistently outperforms sophisticated Transformer-based models across nine real-world datasets, highlighting the need to rethink research directions in time series analysis. We download the official code from: https://github.com/cure-lab/LTSF-Linear for long-term forecasting.
C.3 Implementation Details